Tyler Hawes and Piper Albright formed a partnership, investing $223,200 and $148,800, respectively. Determine their participation in the year's net income of $402,000 under each of the following independent assumptions: No agreement concerning division of net income. Divided in the ratio of original capital investment. Interest at the rate of 15% allowed on original investments and the remainder divided in the ratio of 2:3. Salary allowances of $58,000 and $80,000, respectively, and the balance divided equally. Allowance of interest at the rate of 15% on original investments, salary allowances of $58,000 and $80,000, respectively, and the remainder divided equally.
Answer #1

Determine the share of each partner's in the year's Net Income ( Separately for Individual assumptions )

Contributed Partner's Capital :

Tyler Hawes = $223200

Piper Albright = $148800

Share of Partner's in the Ratio ;

Tyler Hawes : Piper Albright = 223200 : 148800 = 3 : 2

Current Year"s Net Income = $402000

If Interest is Provided at the rate of 15% on Original Investments :

Interest to Tyler Hawes = 223200*15% = $33480

Interest to Piper Albright = 148800*15% = $22320

The following are Independent Assumptions :

1.No agreement concerning division of Net Income :

If there were No Partnership Agreement exists,then the Share of Net Income will be divided equally (1 : 1) by partner"s .

Share of Net Income to Tyler Hawes = 402000*1/2 = $201000

Share of Net Income to Piper Albright = 402000*1/2 = $201000

2.Divided in the ratio of original Investment :

Tyler Hawes : Piper Albright = 223200 : 148800 = 3 : 2

Then, Share of Net Income to Tyler Hawes = 402000*3/5 = $241200

Share of Net Income to Piper Albright = 402000*2/5 = $160800

3. Interest at the rate of 15% allowed on original investments and the remainder divided in the ratio of 2:3 :

Interest is Provided at the rate of 15% on Original Investments :

Interest to Tyler Hawes = 223200*15% = $33480

Interest to Piper Albright = 148800*15% = $22320

Net Income after deduction of Interest :

Net Income     = $402000

Less : Interest = $55800 ($33480+$22320)

  

Net Income after interest $346200

      

Diveded in the ratio of 2:3   :

Share of Net Income after Interest to Tyler Hawes = 346200*2/5 = $138480

Share of Net Income after Interest to Piper Albright = 346200*3/5 = $207720

Total Share of Net Income to Tyler Hawes = $171960 ( $33480 + $138480)

Total Share of Net Income to Piper Albright = $230040 ( $22320 + $207720)

4. Salary allowances of $58,000 and $80,000, respectively, and the balance divided equally.

   Net Income   =    $ 402000

Less : Salary Allowance to Tyler Hawes    = $58000

    Salary Allowance to Piper Albright = $80000

  

Net after Salary Allowance        $264000

  

Balance Net Income After Salary Allowance Divided by Equally

Share of Net Income after Salary Allowance to Tyler Hawes = 264000*1/2 = $132000

Share of Net Income after Salary Allowance to Piper Albright = 264000*1/2 = $132000

Total Share of Net Income to Tyler Hawes = $190000   ( $58000 + $132000)

Total Share of Net Income to Piper Albright = $212000 ( $80000 + $132000)

5. Allowance of interest at the rate of 15% on original investments, salary allowances of $58,000 and $80,000, respectively, and the remainder divided equally.

Interest is Provided at the rate of 15% on Original Investments :

Interest to Tyler Hawes = 223200*15% = $33480

Interest to Piper Albright = 148800*15% = $22320

Net Income After Interest & Salary Allowance :

   Net Income     = $402000

  Less : Interest     = $55800 ($33480+$22320)

      Less : Salary Allowance to Tyler Hawes     = $58000

    Salary Allowance to Piper Albright = $80000

        

Net Income after Interest &Salary Allowance    $208200

  

Balance Net Income After Interest &Salary Allowance Divided by Equally

Share of Net Income after Interest & Salary Allowance to Tyler Hawes = 208200*1/2 = $104100

Share of Net Income after Interest & Salary Allowance to Piper Albright = 208200*1/2 = $104100

Total Share of Net Income to Tyler Hawes = $195580   ($33480+ $58000 + $104100)

Total Share of Net Income to Piper Albright = $206420 ($ 22320+$80000 + $104100)

Thank You !